# HoPE: A Novel Positional Encoding Without Long-Term Decay for Enhanced Context Awareness and Extrapolation

### v1

@@ Abstract
Many positional encodings (PEs) are designed to exhibit long-term decay, based on an entrenched and long-standing inductive opinion: tokens farther away from the current position carry less relevant information. We argue that long-term decay is outdated in the era of LLMs, as LLMs are now applied to tasks demanding precise retrieval of in-context information from arbitrary positions. Firstly, we present empirical analyses on various PEs, demonstrating that models inherently learn attention with only a local-decay pattern while forming a U-shape pattern globally, contradicting the principle of long-term decay. Furthermore, we conduct a detailed analysis of rotary position encoding (RoPE, a prevalent relative positional encoding in LLMs), and found that the U-shape attention is caused by some learned components, which are also the key factor limiting RoPE’s expressiveness and extrapolation. Inspired by these insights, we propose High-frequency rotary Position Encoding (HoPE). HoPE replaces the specific components in RoPE with position-independent ones, retaining only high-frequency signals, which also breaks the principle of long-term decay in theory. HoPE achieves two major advantages: (1) Without constraints imposed by long-term decay, contradictory factors that limit spontaneous attention optimization and model extrapolation performance are removed. (2) Components representing positions and semantics are are optimized. These enhances model’s context awareness and extrapolation, as validated by extensive experiments.
@@ 1 Introduction
However, in the era of LLMs, a question arises: is it still necessary to retain this design? As LLMs are increasingly being applied to long-text and even ultra-long-text scenarios, the importance of distant information appears to depend more on semantics rather than positions.
In this paper, we demonstrate that the answer to the above question is no. Through empirical analyses on various PEs, we found that the attention patterns learned by models tend to contradict the principle of long-term decay. Specifically, models only retain a local-decay pattern, while learning a U-shape attention distribution globally. We further delve into a analysis of RoPE (Rotary Position Encoding by Su et al., 2021, a widely-used RPE), which claims to ensure the long-term decay by combining various frequency components (See Section 4.2.1 for details) while empirically learns the U-shape pattern. We decomposed these components and obtain the following observations:
@@
(2) We explored the attention patterns in extrapolation tasks and found that “activated” components are a key factor limiting RoPE’s extrapolation abilities. These components cause out-of-distribution (OOD) attention logits in the first layer during extrapolation, leading to disarray in subsequent layers’ attention.
(3) The top low-frequency components (whose frequency lower than the “activated” components) tend to stabilize as constant patterns, with a small magnitude. This indicates that these components are not being effectively utilized for representing positional information and learn more about semantics information.
Based on the findings above, we summarize three key insights: (1) Global long-term decay is not necessary for the model and may even hinder optimal learning. (2) To enhance the model’s context awareness and extrapolation, the frequencies of RoPE’s learned components should be constrained. (3) There is redundancy in RoPE, representation subspaces occupied by certain components could be better utilized.
In this paper, we propose a novel positional encoding method called High-frequency rotary Position Encoding (HoPE). HoPE follows above insights and is quite intuitive to implement: we replace the “to-be-activated” and top low-frequency components in the original RoPE with position-independent ones, while retaining the high-frequency components. As a result, contradictory factors for attention optimization are eliminated, extrapolation limitations are reduced, and position information is still well-represented by high-frequency signals.
We conducted extensive evaluations on the model’s extrapolation capabilities and context awareness, including language modeling perplexity, copying tasks, and few-shot learning tasks, both within and beyond the context length. HoPE demonstrates the superior performance compared to other PEs.
@@ 2.2 Relative positional encoding (RPE)
Rotary position encoding (RoPE)
Su et al. (2021) proposed Rotary Position Encoding (RoPE), which encodes positional information by rotating the query and key vectors. The detailed implementation can be found in Section 4.2.1. This encoding method cleverly computes the inner product of relative positions by encoding absolute positions without altering the attention computation process, making it more compatible with various efficient inference methods. However, the original RoPE encoding exhibits poor extrapolation capability for longer sequences Press et al. (2021); Kazemnejad et al. (2023). This raises one popular research direction for exploring RoPE-based length extrapolation methods, such as PI Chen et al. (2023), LongRoPE Ding et al. (2024), Randomized RoPE Ruoss et al. (2023) and YaRN Peng et al. (2023).
Additive relative positional encoding (ARPE)
Another popular method is additive relative positional encoding (ARPE), which introduces a bias matrix $B$ to the original (pre-softmax) attention logits. This approach can be uniformly formula as follows.
${Attn}_{ARPE}(X)=XW_{Q}(XW_{K})^{T}+B$ (1)
Different designs of the bias matrix $B$ result in various APE variants, including T5’s Bias Raffel et al. (2019b), ALiBi Press et al. (2021), KERPLE Chi et al. (2022a), Sandwich Chi et al. (2022b), and FIRE Li et al. (2024). These ARPE methods claim robust performance in length extrapolation without requiring additional operations, as measured by the perplexity (PPL). Nevertheless, some studies Press et al. (2021) noted that PPL may not accurately represent real task performance. Our study further confirms that some ARPEs fail to effectively leverage global information, resulting in only marginal improvements in actual length extrapolation.
3 General Experiments Setups
We train the model using the Llama architecture Touvron et al. (2023a, b) with the task of next-token prediction. The training dataset contains 200 billion tokens sourced from RedPajama Computer (2023). All experiments use Llama tokenizer with a vocabulary of 32,000 tokens and a maximum of 50,000 update steps. Unless otherwise specified, the training length is 512, the model size is 125M and the tests use the last checkpoint. Detailed configurations and other hyperparameters are provided in the Appendix A.
4 Discussion on Long-term Decay in Attention Pattern
In this section, we first present the position-related attention patterns (within the training length) learned by three PEs. We observed that, although the long-term decay of PEs is intuitive, this decay is not global in the empirical attention patterns. Instead, the attention patterns tend to resemble a U-shape curve.
4.1 Attention Patterns based on various PEs
Set up
To observe the position-related attention patterns, we generate 5,000 data samples, each assigning a random token from the vocabulary to all positions in the input sequence (except for the initial [bos]). We then compute the (pre-softmax) attention logit for each position and averaged the results across all heads, noting that most heads exhibited similar patterns. We focused on three PEs including learnable APE, RoPE, and KERPLE. We don’t take ALiBi into account, as its bias matrix $B$ is unlearnable and forces the attention pattern to be global long-term decay.
Result
Results shown in Figure 1. One important observation from the figure is that the attention patterns do not exhibit global long-term decay. Instead, the attention patterns tend to form a U-shape curve, which ensures the decay of adjacent tokens while increasing the importance of the initial tokens.
Moreover, we noted that while RoPE claims it employs multiple components with different frequencies to ensure long-term decay attention, it empirically learns the U-shape pattern. We wonder which components truly matter in this process and delve into a detailed analysis.
Figure 2: (a) Variance Accounted For (VAF) for different components of RoPE during the training phase. (b) The attention pattern of RoPE. The upper subplot displays the attention patterns for each component, while the lower one presents the overall attention pattern along with the combined attention pattern of several “activated” components. (c) The OOD phenomenon in extrapolation caused by “activated” components. The two upper subplots show the attention patterns of the first layer, and the lower subplot presents the anomalous patterns of the subsequent layers. The model training length presented here is 512, results for training length in 1,024 can be found in Appendix B.
4.2 Effects of Different Components of RoPE in Attention Pattern
4.2.1 Preliminaries
Rotary Positional Encoding (RoPE,  Su et al., 2021)
In each Transformer layer, RoPE applies a $d$-dimensional rotation matrix (denoted as $R_{\theta,m}$) to the query or key vector at position $m$ in the sequence for positional encoding. The specific inner product process can be illustrated as follows:
@@ 4.2.1 Preliminaries
$q_{m}\cdot k_{n}$ $=(R_{\Theta,m}q)^{T}(R_{\Theta,n}k)=q^{T}R_{\Theta,m-n}k$
@@
$R_{\Theta,m}=Diag(R_{\theta_{0},m},...,R_{\theta_{d/2-1},m})$
@@
According to the formula above, we can see that the dot product in attention can be broken down into an inner product process of d/2 components, each with a distinct angle $\theta_{i}$, followed by a summation. This can be expressed by the following formula, which allows us to explore the individual effect of each positional component.
$q_{m}\cdot k_{n}=q^{T}R_{\Theta,m-n}k=\sum_{i=0}^{d/2-1}{q_{i}^{T}R_{\theta_{i},m-n}k_{i}}$ (4)
@@
where $i$ represents the component ID, ranging from 00 to $d/2-1$. As shown in the above formula, each component follows the cosine property, with a frequency $\theta_{i}$. And the overall effect of all components theoretically follows the principle of long-term decay(See Section 3.4.3 in Su et al., 2021).
Variance Accounted For (VAF,  Yoon et al., 2021; Qiu et al., 2021)
VAF is primarily used to measure the explanatory power of components for the total variability. It serves as an crucial criterion for identifying effective principal components. A larger value indicates that the component holds greater importance. The formula is as follows:
@@
4.2.2 Experiments
Set up
We adopt a setup similar to that in Section 4.1. Additionally, We calculate the inner product of each component assigned with a unique frequency $\theta$. We then calculate the Variance Accounted For (VAF) every 2,000 training steps to capture key components. Experiments are conducted with training lengths 512 and 1,024.
Result
The results are presented in Figure 2 and Appendix B, from which we can derive three key insights.
(1) The learning of attention patterns is closely associated with some specific components in RoPE, while the model tends to counteract these “activated” components during training. As seen in Figure 2(a) and  6, some components (referred to as “activated” components, i.e. components with IDs 17 to 20 in Figure 2(a)) exhibit high VAF, indicating that they dominate the formation of the overall U-shape pattern. The lower subplot in Figure 2(b) and  7 further confirm this, as the combined pattern of these components mirrors the fluctuations of the overall pattern. However, as indicated in Figure 2(a), the VAF values of the “activated” components decrease as training progresses, suggesting that the model is reducing the contribution of these components. We consider this phenomenon a form of shortcut learning Geirhos et al. (2020); Robinson et al. (2021); Du et al. (2022), which may constrain the model’s overall learning. We also found that all these “activated” components exhibit U-shaped fluctuations across varying training lengths (as seen in the upper subplot of Figure 2(b) and  7). Upon further examination of these components, we found that their frequencies ($\theta$s) fall within the range $\left(\frac{\pi}{L},\frac{2\pi}{L}\right)$, where $L$ is the training length.
@@ 4.2.2 Experiments
(3) Components with a lower frequency than the “activated” ones tend to learn a constant pattern and are not effectively utilized. Another observation is that many components exhibit a constant pattern despite their cosine properties, as shown in the upper subplot of Figure 2(b). Upon delving deeper into these components, we found that their frequencies are all lower than the “activated” components. We speculate that these top low-frequency components do not represent positional information, but rather semantic information. And the properties constraints on them may even hinder this learning, as the corresponding patterns exhibit small magnitudes.
@@ 5 A Novel PE Enhances Model’s Context Awareness and Exploration
Inspired by all experimental observations above, we proposed High-frequency rotary Position Encoding (HoPE). With slight modification in RoPE, HoPE greatly improves the model’s context awareness and extrapolation. We first detail our approach and then validate its effectiveness on perplexity, copy task, and few-shot tasks. The results demonstrate that HoPE exhibits superior performance compared to other PEs.
@@ 5.1 Method
We propose our method based on the following considerations: (1) Global decay is unnecessary, thus some components in position encoding could be removed. (2) Components with U-shape fluctuations within the training length lead to shortcut learning and poor extrapolation. (3) Components with lower frequencies tend to learn semantics but are not be well learned. Since both types of components belong to the low-frequency and are mostly controlled by the latter part of the $R_{\Theta,m}$ matrix in the original RoPE, we implement our approach by replacing these components with position-independent ones while retaining the high-frequency components. We call our method High-frequency rotary Position Encoding (HoPE).
We first identify the “to-be-activated” components and top low-frequency components in original RoPE. As mention in Section 4.2.2, the frequencies ($\Theta_{al}$) and the minimum ID ($a$) of these components could be calculated based on the training context length $L$. The process is as follows:
@@
$a=argmax(\Theta_{al})$
Next, we divide the query (or key) into two parts based on the ID $a$, applying positional encoding only to the first part. For the $R_{\Theta_{h},m}$ matrix applied in positional encoding, we obtain it by setting $\Theta_{h}=\Theta-\Theta_{al}$. The entire process is shown in the following formula.
$q_{m,h}=q_{m}[:2a],q_{l}=q_{m}[2a:]$ (7)
$k_{m,h}=k_{m}[:2a],k_{l}=k_{m}[2a:]$
@@
$R_{\Theta_{h},m}=Diag(R_{\theta_{0},m},...,R_{\theta_{a-1},m})$
$q_{m}\cdot k_{n}=q_{h}^{T}R_{\Theta_{h},m-n}k_{h}+q_{l}^{T}k_{l}$

### v2

@@ Abstract
Many positional encodings (PEs) are designed to exhibit long-term decay, based on an entrenched and long-standing inductive opinion: tokens farther away from the current position carry less relevant information. We argue that long-term decay is outdated in the era of LLMs, as LLMs are now applied to tasks demanding precise retrieval of in-context information from arbitrary positions. Firstly, we present empirical analyses on various PEs, demonstrating that models inherently learn attention with only a local-decay pattern while forming a U-shape pattern globally, contradicting the principle of long-term decay. Furthermore, we conduct a detailed analysis of rotary position encoding (RoPE, a prevalent relative positional encoding in LLMs), and found that the U-shape attention is caused by some learned components, which are also the key factor limiting RoPE’s expressiveness and extrapolation. Inspired by these insights, we propose High-frequency rotary Position Encoding (HoPE). HoPE replaces the specific components in RoPE with position-independent ones, retaining only high-frequency signals, which also breaks the principle of long-term decay in theory. HoPE achieves two major advantages: (1) Without constraints imposed by long-term decay, contradictory factors that limit attention optimization are removed. Thus, the model’s context awareness is enhanced. (2) HoPE exhibits greater robustness to the out-of-distribution behavior in attention patterns during extrapolation. The effectiveness of HoPE is validated through extensive experiments and with a large language model of up to 3 billion parameters.
@@ 1 Introduction
Benchmark RoPE HoPE (Ours)
MMLU 34.27 38.38
MMLU-PRO 12.60 12.74
GPQA 23.23 28.28
BBH 29.00 29.15
WinoGrande 51.70 50.43
GSM8k 10.61 12.05
MATH 1.16 1.84
DROP 31.29 38.46
AVG. 24.23 26.42
Table 1: Performance comparison between RoPE and HoPE across eight benchmarks using a 3B Llama-based model trained with a sequence length of 8192 over 500 billion tokens. Better results are highlighted in bold fonts. HoPE demonstrates superior performance in most tasks.
@@
However, in the era of LLMs, a question arises: is it still necessary to retain this design? As LLMs are increasingly used in long-text scenarios—where they must leverage distant information, such as in Retrieval-Augmented Generation (RAG, Izacard and Grave, 2020)—long-term decay potentially limits their performance.
In this paper, we demonstrate that the answer to the above question is no. Through empirical analyses on various PEs, we found that the attention patterns learned by models tend to contradict the principle of long-term decay. Specifically, models only retain a local-decay pattern, while learning a U-shape attention distribution globally. We further delve into an analysis of RoPE (Rotary Position Encoding by Su et al., 2021, a widely-used RPE), which claims to ensure the long-term decay by combining various frequency components (See Section 2 for details) while empirically learns the U-shape pattern. We decomposed these components and obtained the following observations:
@@
(2) We explored the attention patterns in extrapolation tasks and found that “activated” components are a key factor limiting RoPE’s extrapolation abilities. These components cause out-of-distribution (OOD) attention logits in the first layer during extrapolation, triggering a cascade of disrupted attention patterns through the subsequent layers.
(3) The top low-frequency components (whose frequency is lower than the “activated” components) tend to stabilize as constant patterns, with a small magnitude. This indicates that these components are not being effectively utilized for representing positional information and learn more about semantics information.
Based on the findings above, we summarize three key insights: (1) Global long-term decay is not necessary for the model and may even hinder optimal learning. (2) To enhance the model’s context awareness and extrapolation, the frequencies of RoPE’s learned components should be constrained. (3) There is redundancy in RoPE, and representation subspaces occupied by certain components could be better utilized.
In this paper, we propose a novel positional encoding method called High-frequency rotary Position Encoding (HoPE). HoPE follows above insights and is quite intuitive to implement: we replace the “doomed-to-be-activated” and top low-frequency components in the original RoPE with position-independent ones, while retaining the high-frequency components. As a result, contradictory factors for attention optimization are eliminated, extrapolation limitations are reduced, and position information is still well-represented by high-frequency signals.
On small language models with 125 million parameters, we assess the model’s potential both within and beyond the context length by evaluating perplexity, in-context copying ability and few-shot following ability. HoPE demonstrates superior performance compared to other PEs. We further trained large language models with 3 billion parameters from scratch, and found HoPE performs better than RoPE in complex NLP tasks.
@@ 2.2 Relative positional encoding (RPE)
Rotary position encoding (RoPE, Su et al., 2021) encodes positional information by rotating the query and key vectors. In each Transformer layer, RoPE applies a $d$-dimensional rotation matrix (denoted as $R_{\theta,m}$) to the query or key vector at position $m$ in the sequence for positional encoding. The specific inner product process can be illustrated as follows:
@@
$q_{m}\cdot k_{n}$ $=(R_{\Theta,m}q)^{\top}(R_{\Theta,n}k)=q^{\top}R_{\Theta,m-n}k$
@@
$R_{\Theta,m}=\texttt{Diag}(R_{\theta_{0},m},...,R_{\theta_{d/2-1},m})$
@@
This encoding method cleverly computes the inner product of relative positions by encoding absolute positions without altering the attention computation process, making it more compatible with various efficient inference methods. However, the original RoPE encoding exhibits poor extrapolation capability for longer sequences Press et al. (2021); Kazemnejad et al. (2023). This raises one popular research direction for exploring RoPE-based length extrapolation methods, such as PI Chen et al. (2023), LongRoPE Ding et al. (2024), Randomized RoPE Ruoss et al. (2023) and YaRN Peng et al. (2023).
Additive relative positional encoding (ARPE) is another popular method, which introduces a bias matrix $B$ to the original (pre-softmax) attention logits. This approach can be uniformly formula as follows.
${Attn}_{ARPE}(X)=XW_{Q}(XW_{K})^{T}+B$ (3)
Different designs of the bias matrix $B$ result in various APE variants, including T5’s Bias Raffel et al. (2019b), ALiBi Press et al. (2021), KERPLE Chi et al. (2022a), Sandwich Chi et al. (2022b), and FIRE Li et al. (2024). These ARPE methods claim robust performance in length extrapolation, as measured by the perplexity (PPL). Nevertheless, some studies Press et al. (2021) noted that PPL may not accurately represent real task performance. Our study further confirms that some ARPEs fail to effectively leverage global information, resulting in only marginal improvements in actual length extrapolation.
3 Discussion on Position-related Attention Pattern
In this section, We first present the position-related attention patterns (within the training context length) learned by three PEs. We observed that, although the long-term decay of PEs is intuitive, this decay is not global in the empirical attention patterns. Instead, the attention patterns tend to resemble a U-shape curve.
3.1 Experiments Setups
We train small Llama language models Touvron et al. (2023a, b); Dubey et al. (2024) with 125 million parameters, using different PEs. The training dataset contains 200 billion tokens sourced from RedPajama Computer (2023). All experiments use Llama tokenizer with a vocabulary of 32,000 tokens. The training context length is 512 tokens and the update steps are 50,000. Detailed configurations and other hyperparameters are provided in the Appendix A.
To observe the position-related attention patterns both within and beyond the training context length, we set two test lengths: 512 and 1024. For each test length, we generate 5,000 corresponding data samples, each assigning a random token from the vocabulary to all positions in the input sequence (except for the initial [bos]). We then calculate the pre-softmax attention logits for each position. To illustrate a common pattern, we average the results across all layers and heads, as most heads demonstrate similar behaviors.
3.2 Long-term Decay in Attention Patterns
Our analysis focused on three PEs including learnable APE, RoPE, and KERPLE. We don’t take ALiBi into account, as its bias matrix $B$ is unlearnable and forces the attention pattern to be global long-term decay.
Results are shown in Figure 1. One important observation is that the attention patterns do not exhibit global long-term decay. Instead, the attention patterns tend to form a U-shape curve, which ensures the decay of adjacent tokens while increasing the importance of the initial tokens.
Figure 2: (a) We decomposed RoPE into components (Comps.) for analysis (See Eq.4). The upper subplot displays the contribution to the overall attention logits from each component. We highlight some components with outstanding patterns, namely “activated” components, in red and lower frequency components in blue. The lower subplot presents the overall attention logits, along with the combined effects of “activated” components we highlighted in the top figure. (b) Variance Accounted For (VAF, See Section 3.3.1) for different components of RoPE during training. (c) The OOD phenomenon in extrapolation caused by “activated” components. The two upper subplots show the attention patterns of the first layer, and the lower subplot presents the anomalous patterns of the subsequent layers. The model training length presented here is 512, results for training length in 1,024 can be found in Appendix B.
3.3 Effects of Different Components of RoPE in Attention Pattern
Figure 1 also demonstrated that RoPE empirically learns the U-shape pattern while claiming to employ multiple components with different frequencies to ensure long-term decay attention. We wonder which components truly matter in this process and delve into a detailed analysis.
3.3.1 Preliminaries
Components of RoPE
According to Eq.2, we can see that the dot product in attention can be broken down into an inner product process of d/2 components, each with a distinct angle $\theta_{i}$, followed by a summation. This can be expressed by the following formula, which allows us to explore the individual effect of each positional component $C_{i}$.
$q_{m}\cdot k_{n}=q^{T}R_{\Theta,m-n}k=\sum_{i=0}^{d/2-1}\underbrace{{q_{i}^{T}R_{\theta_{i},m-n}k_{i}}}_{C_{i}}$ (4)
@@ 3.3.1 Preliminaries
Variance Accounted For (VAF)
VAF Yoon et al. (2021); Qiu et al. (2021) is primarily used to measure the explanatory power of components for the total variability. It serves as a crucial criterion for identifying effective principal components. A larger value indicates that the component holds greater importance. The formula is as follows:
@@
3.3.2 Experiments and Results
We decompose RoPE into several components, each associated with a unique frequency $\theta$, and examine their individual contribution to the overall attention logits, both within the training length and during extrapolation. Additionally, under the scenarios within training context length, we tracked the pattern dynamics of components as training progresses, using VAF metric.
The results are presented in Figure 2.111To further validate our findings, we also included results with a longer training context length of 1,024 in Appendix B. From these results, we can derive three key insights.
(1) The learning of attention patterns is closely associated with some specific components in RoPE, while the model tends to counteract these “activated” components during training. As seen in Figure 2(b), some components (referred to as “activated” components, highlighted with red in Figure 2(b)) exhibit high VAF, indicating that they dominate the formation of the overall U-shape pattern. The lower subplot in Figure 2(a) further confirms this, as the combined pattern of these components mirrors the fluctuations of the overall pattern. However, as indicated in Figure 2(b), the VAF values of the “activated” components decrease as training progresses, suggesting that the model is reducing the contribution of these components. We consider this phenomenon a form of shortcut learning Geirhos et al. (2020); Robinson et al. (2021); Du et al. (2022), which may constrain the model’s overall learning. We also found that all these “activated” components exhibit U-shaped fluctuations across varying training lengths (as seen in the upper subplot of Figure 2(a)). Upon further examination of these components, we found that their frequencies ($\theta$s) fall within the range $\left(\frac{\pi}{L},\frac{2\pi}{L}\right)$, where $L$ is the training length.
@@ 3.3.2 Experiments and Results
(3) Components with a lower frequency than the “activated” ones tend to learn a constant pattern and are not effectively utilized. Another observation is that many components exhibit a constant pattern despite their cosine properties, as shown in the upper subplot of Figure 2(a). Upon delving deeper into these components, we found that their frequencies are all lower than the “activated” components. We speculate that these top low-frequency components do not represent positional information, but rather semantic information. And the properties constraints on them may even hinder this learning, as the corresponding patterns exhibit small magnitudes.
@@ 4 A Novel PE Enhances Model’s Context Awareness and Exploration
Inspired by all experimental results and observations above, we proposed High-frequency rotary Position Encoding (HoPE). With slight modification in RoPE, HoPE greatly improves the model’s context awareness and extrapolation. We first detail our approach and then validate its effectiveness on perplexity, copying task, and few-shot following tasks. The results demonstrate that HoPE exhibits superior performance compared to other PEs.
@@ 4.1 Method
We propose our method based on the following considerations: (1) Global decay is unnecessary, thus some components in position encoding could be removed. (2) Components with U-shape fluctuations within the training length lead to shortcut learning and poor extrapolation. (3) Components with lower frequencies tend to learn semantics but are not well learned. Since both types of components belong to the low-frequency and are mostly controlled by the latter part of the $R_{\Theta,m}$ matrix in the original RoPE, we implement our approach by replacing these components with position-independent ones while retaining the high-frequency components. We call our method High-frequency rotary Position Encoding (HoPE).
We first identify the “doomed-to-be-activated” components and top low-frequency components in the original RoPE. As mentioned in Section 3.3, the frequencies ($\Theta_{al}$) and the minimum index $a$ of these two components could be calculated based on the training context length $L$. The process is as follows:
@@
$a=\texttt{argmax}(\Theta_{al})$
Next, we divide the query (or key) into two parts based on the index $a$, applying positional encoding only to the first part. For the $R_{\Theta_{h},m}$ matrix applied in positional encoding, we obtain it by setting $\Theta_{h}=\Theta-\Theta_{al}$. The entire process is shown in the following formula.
$q_{m,h}=q_{m}[:2a],k_{n,h}=k_{n}[:2a]$ (7)
@@
$R_{\Theta_{h},m}=\texttt{Diag}(R_{\theta_{0},m},...,R_{\theta_{a-1},m})$
$q_{m}\cdot k_{n}=q_{h}^{\top}R_{\Theta_{h},m-n}k_{h}+q_{l}^{\top}k_{l}$